# Visibility-Aware RRT* for Safety-Critical Navigation of Perception-Limited Robots in Unknown Environments

### v1

@@ VII CONCLUSION
In this paper, we considered that safety may not be guaranteed for robots with limited perception capabilities navigating in unknown environments, even when using safety-critical controllers. To address this challenge, we introduced the Visibility-Aware RRT* algorithm, which generates a global reference path that is both collision-free and visibility-aware, enabling the tracking controller to avoid hidden obstacles at all times. We introduced the visibility CBF to encode the robot’s visibility limitations, utilizing its constraint as a termination criterion during the steering process, alongside a standard collision avoidance CBF constraint. We also proved that our algorithm ensures the satisfaction of visibility constraints by the local controller, and we evaluated the performance of the integrated path planner with two different controllers via simulations. The results indicated that our method outperformed all other compared baselines in both safety and efficiency aspects.

### v2

@@ VII CONCLUSION
In this paper, we considered that safety may not be guaranteed for robots with limited perception capabilities navigating in unknown environments, even when using safety-critical controllers. To address this challenge, we introduced the Visibility-Aware RRT* algorithm, which generates a global reference path that is both collision-free and visibility-aware, enabling the tracking controller to avoid hidden obstacles. We introduced the visibility CBF to encode the robot’s visibility limitations, utilizing its constraint as a termination criterion during the steering process, alongside a standard collision avoidance CBF constraint. We also proved that our algorithm ensures the satisfaction of visibility constraints by the local controller, and we evaluated the performance of the integrated path planner with two different controllers via simulations. Our method outperformed all other compared baselines in both safety and efficiency aspects.